OpenStudy (anonymous):

How do you solve 3+10^2x=8

OpenStudy (anonymous):

Well first subtract 3 from both sides 10^2x=5 then take log base 10 of both sides log(10^2x)=log(5) left cancels 2x=log(5) divide by 2 x=log(5)/2